OpenStudy (anonymous):

t(7t+4)=0 (fractions and and itegers)

OpenStudy (anonymous):

t(7t+4)=0 7t^2+4t = 0 a = 7, b = 4 , c=0; t = (-b+- sqrt(b^2 - 4.a.c))/2a t= (-4+- sqrt(16 -0))/14 t = (-4 +- 4)/14 for (+) t =(-4+4)/14 = 0 for (-) t=(-4-4)/14 = -8/14 = -4/7 so t =0 or -4/7